### v1

# Channel Twinning: An Enabler for Next-Generation Ubiquitous Wireless Connectivity

Abstract
The emerging concept of channel twinning (CT) has great potential to become a key enabler of ubiquitous connectivity in next-generation (xG) wireless systems. By fusing multimodal sensor data, CT advocates a high-fidelity and low-overhead channel acquisition paradigm, which is promising to provide accurate channel prediction in cross-domain and high-mobility scenarios of ubiquitous xG networks. However, the current literature lacks a universal CT architecture to address the challenges of heterogeneous scenarios, data, and resources in xG networks, which hinders the widespread deployment and applications of CT. This article discusses a new modularized CT architecture to bridge the barriers to scene recognition, cooperative sensing, and decentralized training. Based on the modularized design of CT, universal channel modeling, multimodal cooperative sensing, and lightweight twin modeling are described. Moreover, this article provides a concise definition, technical features, and case studies of CT, followed by potential applications of CT-empowered ubiquitous connectivity and some issues requiring future investigations.
Index Terms:
Next-generation (xG) wireless communication system, channel twinning (CT), ubiquitous connectivity, multimodal sensing, lightweight digital twin.
I Introduction
Ubiquitous connectivity is regarded as an overarching and fundamental function of next-generation (xG) networks (e.g., 6G networks), as it underpins seamless wireless coverage and network resilience to natural disasters [1]. To have ubiquitous connectivity, advanced coverage extension solutions such as space-air-ground-sea-integrated network (SAGSIN), cell-free network, and reconfigurable intelligent surface (RIS) have been proposed [2]. The provisioning of zero coverage-hole networks has to deal with the challenges related to diversified wireless channels across heterogeneous networks and multiple frequency bands, thus rendering network planning and optimization more convoluted than ever. This necessitates attention to one of the fundamental challenges in wireless communication systems, namely, channel acquisition, which determines the practical performance limits of wireless systems and networks.
I-A Paradigm Shifts in Channel Acquisition
Recently, channel twinning (CT) has emerged as a new regime of channel acquisition that reflects the complex interactions of channels and physical environments. CT has received increasing attention because it subverts traditional channel acquisition in the following research paradigms.
I-A1 A shift of perspective from mathematical to environmental modeling
Mathematical models (statistical and deterministic models) cannot balance the accuracy and complexity [3]. Instead, CT aims to reveal the channel mapping from environmental attributes, and achieve an accurate channel acquisition with low complexity.
I-A2 A shift from unimodal to multimodal sensing
Traditional channel acquisition primarily relies on radio frequency (RF) sensing-based pilot training methods, accompanied by high overhead. In contrast, multimodal fusion is a pillar of CT, and it is promising to reduce or even eliminate RF pilots [4].
I-A3 A shift from physical-world to digital-world measurement
In xG networks, cross-domain and high-mobility environments can cause poor channel estimation accuracy. While channel prediction can apply to dynamic scenarios, it is often not environment-aware and is limited by the correlation model. By emulating the interactions between the channel and the physical world, CT can create real-time channel replicas in the digital world.
Table I provides a conceptual comparison to clarify the relationship between CT and other related concepts.
TABLE I: Comparison between different concepts about channel acquisition.
Concept
Role
Execution
phase
Modeling
Method
Accuracy
Generalizability
Complexity
Environment
synchronization
Research
perspective
Chanel modeling Derive proper models for effective channel representation Before network deployment
Statistical
modeling
Low $\times$× Low $\times$× Mathematical representation
Deterministic
modeling
High $\times$× High $\times$×
Hybrid
modeling
Moderate $\times$× Moderate $\times$×
Channel
estimation
Obtain current
channel parameters
After network
deployment
Pilot-based or
training-based
Moderate $\times$× Moderate $\times$×
Mathematical
representation
Channel
prediction
Infer future
channel states
After network
deployment
Time series
analysis
Moderate $\times$× Moderate $\times$×
Mathematical
representation
Channel twinning Create the digital replica of the channel evolution Full life cycle Synthetic CT High $\times$× Low $\times$× Environment semantics
Transfer CT Moderate Moderate
Multiverse CT Multi-level High
I-B Current Status of CT Studies
In the context of CT, some initial studies have been reported in [5, 6, 7, 8]. The term of DT channel (DTC) [^1] To emphasize that “channel twinning” is a technology, similar to channel modeling or channel sounding, we have not used the terminology “DTC”. appeared earlier in [5], but the concept has not been specified. Later, the potential applications of predicting channels via digital twin (DT) were highlighted in [6], although the DTC concept was not proposed. The authors of [7] mentioned that DTC should enable channel prediction but did not provide clear definitions. More recently, DTC has been defined as the mapping between channels and associated communication operations [8]. Therein, the feasibility of real-time DTC has been demonstrated experimentally. As an initial instantiation of the CT concept, Boston Twin [9] was developed to introduce a city-scale twinning system for wireless networks.
The above literature has focused primarily on beam prediction applications or only discusses the knowledge graph construction from a local and centralized perspective. However, deploying CT in xG networks faces the challenges of cross-domain modeling, distributed sensing, and lightweight training. A centralized strategy makes it hard to cater to heterogeneous resources in xG networks.
Therefore, in this article, a systematic discussion on the modularized architecture of ubiquitous CT is conducted to facilitate the organic integration of CT and ubiquitous sensing and intelligence. For a better understanding of CT, a concise definition and several clarifications regarding the general CT misconceptions are refined. This article also highlights the technical evolutions, the potential use cases for CT-empowered ubiquitous connectivity technologies, and several promising directions to spur future research.
Fig. 1: Conceptual framework of CT in xG networks with ubiquitous connectivity. Advanced ubiquitous connectivity scenarios give birth to diversified wireless channels across heterogeneous networks. All of these channels can be acquired in the CT system without the need for high-overhead channel estimation.
II Basic Concepts and Features of CT
II-A Concepts and Realization Principle of CT
In this article, CT is defined as a digital mirror construction of physical channels for network design, optimization, and evaluation without real deployment. Essentially, such a digital mirror is a mapping model that indicates the correspondence from the environment to the associated channel state information (CSI). Specifically, a site-specific CT attempts to find a mapper to approximate the RF propagation law. As shown in Fig. 1, given the environmental conditions, CT can acquire the CSI at any position.
Regarding the realization of CT, current mainstream methods can be primarily divided into three categories: Spatial interpolation modeling, environment semantic modeling, and deep learning (DL) modeling.
II-A1 Spatial Interpolation Modeling
Spatial interpolation [2] only utilizes channel measurements within the region of interest (RoI) to predict the channel of any location.
II-A2 Environment Semantic Modeling
Environment semantic-based methods, such as spatial loss field and virtual obstacle model [10], estimate the fading distribution of the entire RoI by observations.
II-A3 Deep Learning Modeling
In DL methods, differentiable ray tracing (DRT) is gaining popularity as a promising DT enabler [11]. The DRT principle is to learn the electromagnetic (EM) properties of involved materials by modeling the interaction between ray paths and the environment.
In practice, suitable methods should be selected according to applications and dataset sizes. From the perspective of implementation complexity, CT deployment should undergo three evolution stages: i) Synthetic CT [7], which is deployed in a quasi-static environment to generate synthetic data that approximates the measurements in the RoI; ii) Transfer CT [12], which constructs self-adaptive CT based on knowledge transfer to cope with unforeseen scenarios; and iii) Multiverse CT [13], which customizes different twins for applications with varying computing power and delay constraints. Each twin model offers multiple fidelity levels with different processing delays.
II-B Common Misconceptions and Three Clarifications
To elaborate on the concept of CT comprehensively, we discuss the CT model illustrated in Fig. 1 from inputs to outputs. Several basic features are highlighted as follows: a) context-awareness: Instead of merely inputting received pilots or feedback by RF sensing methods, CT integrates contextual information from multimodal sensors (e.g., camera, GPS and Radar) to promptly adapt to environmental dynamics; b) high-fidelity: CT allows for high-fidelity emulation of complex interactions between the environment and the network. Thanks to three-dimensional (3D) map-aided methods, e.g., ray-tracing, spatial consistency [1] and multi-hop propagation can be simulated; c) self-synchronization: CT can model the whole life cycle of channel behaviors in the wireless environment for real-time emulation; d) expertise-driven: EM and channel modeling expertise are needed to design CT models. This helps produce compact DL structures and improve interpretability when specific channel characteristics, e.g., signal-to-noise ratio (SNR), are desired to be output; and e) region-specific: Since the CT model is trained on measurements from a specific site’s coverage region, the DTC must be region-dependent.
Since CT is an emerging technology, there are several common misconceptions that need to be clarified.
Clarification 1: Do widely-used statistical channel modeling techniques belong to CT technology?
Following the CT concept, a DTC refers to a digital mirror of the channel response in a given environment, whereas standard statistical models like Rayleigh and Rician, cannot match specific environments. Moreover, 3GPP urban microcell (UMi) and urban macrocell (UMa) models only consider macroscopic environmental attributes (e.g., urban or rural) but ignore the specific environmental factors. In this sense, statistical models may not create the real DTCs.
Clarification 2: Does CT necessarily mean the use of DL?
According to specific application requirements, one can flexibly choose different CT training methods, e.g., spatial interpolation, environment semantic, and DL methods.
Clarification 3: Does using only position information to predict network operations belong to CT technology?
CT is either explicitly or implicitly context-aware. Only position-based methods construct the channels under the simple line-of-sight (LoS) assumption that ignores the environmental factors. However, blockages and scatters may cause non-LoS (NLoS) cases. This shows the significance of context awareness, and that position-based channel recovery cannot compete with CT.
III A Modularized Design of CT Architecture
For the design and deployment of CT in xG networks, the following questions need to be answered.
  • Q1.
    Since xG networks are anticipated to involve massive cross-domain links across space/frequency/time-domains, how can we make the channel modeling of CT applicable to any scenario?
  • Q2.
    How do we effectively integrate multimodal sensor networks during the CT deployment?
  • Q3.
    How can CT training in network nodes with limited computing capacity and power supply be realized?
In light of these questions, we present a modularized architecture to discuss the design considerations for CT deployment in xG networks. The overall CT architecture comprises three modules of universal channel modeling, multimodal cooperative sensing, and lightweight decentralized CT training, as shown in Fig. 2.
Fig. 2: The overall modularized CT architecture. The universal channel modeling module provides information about the sensing targets and prediction tasks of interest for the other two modules. The channel modeling module can generate synthetic data, which can be leveraged to pretrain the CT model in the lightweight twin module. Then, the CT model can be fine-tuned by feeding sensing measurements from the multimodal cooperative sensing module.
III-A Universal Channel Modeling Module
Various scenarios in xG networks give rise to various types of channels. The characteristics of these channels differ significantly from each other. Consequently, a major challenge that arises for CT deployment is how to integrate the diverse channel characteristics across scenarios into a unified model. To this end, a universal channel modeling module tailored for ubiquitous CT is needed.
Generally, a channel consists of large-scale fading (LSF) and small-scale fading (SSF) components, each with different characteristics in various scenarios. Therefore, the channel components associated with the scenario and working band should be identified at first. For instance, high-mobility scenarios, such as Internet of Vehicles (IoVs) and unmanned aerial vehicle (UAV) networks, should consider non-stationary channels. Gas absorption and blocking effects should be considered in millimeter-wave (mmWave) and Terahertz (THz) networks, while visible light communication (VLC) channels only involve LSF. Extremely large-scale, multiple-input-multiple-output (MIMO) channels should consider spherical wavefront and spatial non-stationarity [1].
To effectively deploy CT, we can quickly determine the channel components by following two steps: 1) LSF channel construction by a static 3D map; and 2) SSF channel prediction by capturing dynamic activities.
III-B Multimodal Cooperative Sensing Module
With the rise of Internet of Things (IoT) devices and advances in integrated sensing and communication (ISAC), multimodal sensing has become increasingly common. Multimodal sensing data fusion (M2SDF) is imperative for CT model training [4]. In the multimodal cooperative sensing module shown in Fig. 2, there are two main categories of multimodal sensors: RF and non-RF.
The traditional channel measurement sources come primarily from RF reception and sensing. Such measurement is limited by the sensing range and accuracy of the single-site ISAC scheme. For RF sensing, a multi-site cooperative sensing scheme can be preferable, which can enhance signal acquisition by virtue of ubiquitous connectivity. In this scheme, non-terrestrial and terrestrial base stations (BSs) can sense any position through ISAC technology. In addition, users can also track targets by the dedicated positioning reference signals defined in 5G-NR R17 [14].
However, cooperative sensing requires precise space-time registration, which is crucial for positioning data acquisition and continuous updates of CT systems. Non-RF sensing involves more heterogeneous multimodal data than RF sensing. In this sense, the challenge of M2SDF lies in how to deal with the data heterogeneity at the CT server. To extract propagation environment semantics, semantic alignment and completion [8] techniques may be invoked.
Fig. 3: Training architecture of CT: (a) Centralized learning-based CT model construction, where dedicated measurement anchors collect data and upload data to the CT server for training; and (b) FL-based CT model construction, where registered device clients use their local data to train the model and only upload trained parameters for model aggregation at the CT server.
III-C Lightweight Twin Modeling Module
Existing CT models typically train in a centralized manner. However, scalability and privacy challenges limits the CT deployment in ubiquitous connectivity. In this case, federated learning (FL) is viewed as an adequate and promising choice for lightweight CT because FL training is decentralized and does not require sharing data. Fig. 3 shows a lightweight CT training module based on FL. In the FL-enabled CT scheme, FL clients only need to upload local model parameters to collaboratively train a global CT model, thus reducing communication costs and privacy leakage risks.
The integration of FL and DT is not straightforward, as two practical issues need to be addressed as prerequisites. One is that participating in FL tasks can increase the energy costs of client devices. For this reason, self-interested sensor devices may be reluctant to participate in FL, which needs proper FL incentive mechanisms [15]. Clients participating in FL should meet their own energy constraints. For example, UAVs are energy-limited. In this sense, an incentive mechanism for FL-based CT should be designed to improve CT training efficiency while meeting the power constraints. To achieve such energy-efficient FL incentives, we need to quantify the contribution of clients to the global CT model convergence and design appropriate payment mechanisms. The other practical issue that needs to be addressed for the integration of FL and DT is the heterogeneity of local data on each client. This would require handling data distribution imbalances across clients. The reason is that personalized FL may be suitable for lightweight CT in the future.
IV Case Studies: Intelligent Coverage and Lightweight Twin
This section provides two case studies to demonstrate the significance of CT for ubiquitous connectivity and discuss the lightweight FL-based CT performance.
Fig. 4: User association results produced based on the true channel, the twin channel, and the fitted PL model: (a) True channel gain; (b) twin channel gain; (c) channel gain predicted by fitted PL model; and (d) the achievable sum rate with different schemes. Each UE selects five APs for communications, and the associated APs are indicated by red dots.
Fig. 5: Federated twin of the channel gains: (a) The validation loss versus communication rounds, where the total number of training samples is 10,000, and these samples are randomly divided into three or five UEs for FL; and (b) MAE comparison of channel gain prediction across different models.
IV-A CT-Based Intelligent Coverage
In ubiquitous connectivity networks, user association can dynamically adjust the connection of user equipment (UE) according to the channel status, in order to achieve continuous and efficient coverage. By user association, the optimal access points (APs) are selected for UE access, affecting network resource utilization. Dynamic adaptive user association can reduce latency or packet loss, avoid interruption, and save the transmit powers. However, the premise of dynamic adaptive user association is an accurate perception of full CSI.
Cell-free communication, as a key ubiquitous connectivity technology, is taken as an example to illustrate the advantage of CT. In cell-free networks, obtaining the full CSI of involved nodes is challenged by limited pilots and high CSI exchange overhead. To address this, we propose a CT-based intelligent user association scheme in cell-free networks. Specifically, the proposed scheme is implemented via offline CT model training and online CSI query, as illustrated in Fig. 3(a). In the offline phase, dedicated measurement anchors are responsible for collecting channel measurements at various locations in the RoI and uploading measurements to the CT server for training. Once the CT model is obtained through offline training, it is stored in the twin repository. In the online phase, APs or UEs request the CT server to query the channel information of the node they attempt to communicate with.
To demonstrate the advantage of CT, we leverage the synthetic data generated by Wireless InSite [^2] https://www.remcom.com/wireless-insite-propagation-software software for model training [2]. Channel gain measurements, sampled at a uniform grid in the region shown in Fig. 4(a), serve as the ground truth.
Based on the spatial interpolation concept, the trained CT model can capture the impact of environmental factors on channel gains. Fig. 4(b) shows a channel gain map given by a UE position, where the UE is indicated by a white square. Conventional user association is based only on the distance or path loss (PL) in practical cell-free systems. We derive the PL model with the same synthetic data by least squares fitting. As a benchmark, PL-based results are shown in Fig. 4(c). By comparing Figs. 4(a), 4(b), and 4(c), it can be observed that CT can efficiently perceive the environment, while the fitted PL model cannot reproduce the environment profile.
With the predicted channel gain, we select five APs from 20 distributed APs within the RoI for user association. Since the PL model does not grasp the environment semantics, it can only select APs according to the distance criterion. The channel gain comprises PL, shadowing, and multi-path effects. PL is determined by distance while shadowing and multi-path effect are affected by surroundings.
By contrast, CT can guide the UE in avoiding buildings, instead of directly selecting the nearest APs. Fig. 4(d) shows the sum rates achieved by different user association schemes, which confirms the benefits of CT for ubiquitous coverage. In addition, since open street maps [10] and user locations are reasonably easy to obtain, the CSI of any AP-UE link can be retrieved by CT. Thus, the dreadful CSI-sharing overhead can be avoided in ubiquitous connectivity.
IV-B Federated Learning-Based Lightweight CT
Centralized training of DT relies on sufficient computing capability and energy supply, which may be unavailable in sensor devices, e.g., UAVs. Fortunately, FL can realize a decentralized DT model construction while avoiding data privacy issues. In light of this, we investigate an FL-based lightweight channel gain twinning model; see Fig. 3(b).
In the FL-based CT training architecture, no dedicated measurement anchors are needed for data collection. Each mobile UE may be a potential measurement anchor. These incentive UEs can record their locations and associated channel measurements as they move into the RoI. Unlike centralized training, the UEs do not upload the data; rather, they first register their identities with the CT server’s client manager for subsequent FL training. When the CT server prepares to train a CT model, a group of registered UEs train their own CT models locally by using their recorded measurements. Then, they upload only the trained parameters to the server for model aggregation, thus obtaining the global CT model. The CT server sends the global model parameters to the registered clients to update the local models. After several rounds of interactions, the global model converges, and hence, the FL-based CT training is completed.
We select a portion of the channel gain dataset used in the previous subsection, randomly splitting and distributing it to FL clients. Fig. 5(a) shows the trained results. It is seen that more clients contribute to better validation loss performance, even if the total data samples remain the same. With more clients, the gradient noise of individual clients can be smoothed out. This indicates that increasing the number of clients can improve the federated twin performance, but raises communication overhead concerns.
Fig. 5(b) compares prediction performance among the centralized CT model, fitted PL model, and federated CT model in terms of mean absolute error (MAE). From the boxplots of MAE, the red solid line denotes the median MAE, the box represents the distribution range of the majority of MAE, and the black dots outside the box denote outliers. It is observed that the PL model exhibits high prediction errors overall, indicating poor prediction accuracy. Meanwhile, few outliers indicate prediction stability, as the channel gain produced by the PL model lacks shadowing and SSF compared to the true channel. Both CT models show better prediction accuracy. Even decentralized training-based CT model has superior accuracy to the PL model. Despite the fluctuations in prediction error of CT models, occasional outliers do not affect the overall performance of ubiquitous connectivity. Notably, the goal of CT is not to copy physical channel entities, but rather to satisfy network requirements by emulation.
V Convergence of CT and Ubiquitous Communications: Opportunities and Challenges
V-A Opportunities of CT for Ubiquitous Connectivity
V-A1 CT-Empowered Non-Terrestrial Communications
Due to the high mobility of non-terrestrial systems, both satellite-to-ground and UAV-to-ground channels suffer from the non-negligible channel aging effect. Unlike the ground channel, UAV-to-ground channel gains do not show smooth changes with the movement of the UAVs. Due to the down-tilted antenna pattern, UAVs may easily fall into the antenna nulls of nearby BSs. Therefore, the UAVs must search for the best antenna sidelobe associated with the remote cell when switching cells instead of using the aforementioned distance criterion. In this case, CT can emulate the antenna gain and airframe shadowing [1] to achieve high-efficiency cell association in UAV networks.
V-A2 CT-Empowered IoV Communications
One of the most immediate applications of CT is predictive beamforming in IoVs. CT-enabled beamforming must only send necessary control signaling without extra pilots [4]. On the other hand, RIS is regarded as an attractive low-power technology that can overcome dynamic blockage in IoVs. However, CSI acquisition and frequent beam alignment in high-mobility scenarios hinder the widespread application of RIS. Through the CT model dedicated to serving the RIS sites, adaptive beam width control and vehicle tracking can be further realized.
V-A3 CT-Empowered Maritime Communications
CT is promising in achieving ubiquitous maritime connectivity. Note that the maritime communications environment has distinct features of sparse scatters compared to terrestrial networks. This motivates us to employ the environmental information to deploy CT models for network planning [3]. Channel interference can be acquired in advance to avoid the acquisition burden of full CSI. In the harsh underwater environment, underwater acoustic (UWA) channels suffer from spatio-temporal-frequency variability, limited bandwidth and severe noise and interference [1]. Channel estimation and data transmission becomes even more challenging in multi-node UWA communications. In this case, CT also brings unprecedented opportunities for UWA communications.
V-B Challenges and Open Issues
V-B1 Data Deficiency
CT demands sufficient high-quality measurement data. However, channel measurement campaigns are time-consuming and expensive. In practice, only sparse measurements from limited points are available for CT training. For example, DL-based indoor propagation prediction [12] only utilizes channel measurements from 36 locations. Furthermore, some measurements may contain errors due to hardware defects. Therefore, for high-precision CT training, it is necessary to study the efficient utilization of synthetic and measurement data.
V-B2 Personalized FL Mechanism of CT
Since FL clients only collect channel measurements from the local irregular areas, a simple federated averaging-based aggregation strategy can hardly be optimal. In particular, personalized FL with incentive mechanisms needs to consider the CT model contribution of each client to the entire environment map.
V-B3 Malicious CT Deployment
Malicious nodes with CT ability may exist in networks. They can train the CT models based on local radio measurements or query the CT server for CSI at locations of interest. Therefore, once a malicious CT is able to recover CSI, wireless networks will face unparalleled security threats. Existing physical layer security and secret key schemes will be rendered ineffective.
V-B4 High-Dimensional DTC Representation
While DTC is effective for wireless network optimization, high-dimensional channel matrices caused by large-scale MIMO or RIS raise computational and caching challenges for CT deployment. In this case, an equivalent channel compression representation scheme in CT is attractive for network optimization.
VI Conclusion
From the perspective of the physical environment, a new channel acquisition paradigm, namely, CT technology, has been studied to support xG ubiquitous wireless communication networks. With multimodal sensing, CT can evolve to predict CSI that closely resembles actual channels. This article outlines recent CT-related activities with different conceptual names. Then, the concept, key features, and two case studies of CT have been discussed. To promote CT deployment in wireless networks, a modularized design architecture of CT has been presented. The architecture includes universal channel modeling of full scenarios, multimodal cooperative sensing, and FL-based lightweight CT training. New opportunities for CT in enabling technologies for ubiquitous connectivity and future research challenges have also been discussed.

### v3

@@
-Channel Twinning: An Enabler for Next-Generation Ubiquitous Wireless Connectivity
+Advancing Ubiquitous Wireless Connectivity through Channel Twinning
 Abstract
-The emerging concept of channel twinning (CT) has great potential to become a key enabler of ubiquitous connectivity in next-generation (xG) wireless systems.
-By fusing multimodal sensor data, CT advocates a high-fidelity and low-overhead channel acquisition paradigm, which is promising to provide accurate channel prediction in cross-domain and high-mobility scenarios of ubiquitous xG networks.
-However, the current literature lacks a universal CT architecture to address the challenges of heterogeneous scenarios, data, and resources in xG networks, which hinders the widespread deployment and applications of CT.
-This article discusses a new modularized CT architecture to bridge the barriers to scene recognition, cooperative sensing, and decentralized training.
-Based on the modularized design of CT, universal channel modeling, multimodal cooperative sensing, and lightweight twin modeling are described.
-Moreover, this article provides a concise definition, technical features, and case studies of CT, followed by potential applications of CT-empowered ubiquitous connectivity and some issues requiring future investigations.
+As an emerging trend in channel acquisition (CA), the concept of channel twinning (CT) has been proposed as a powerful enabler of ubiquitous connectivity in next-generation (xG) wireless systems.
+By fusing multimodal sensor data, CT advocates a high-fidelity and low-overhead CA paradigm, which is promising to provide accurate channel prediction in cross-domain and high-mobility scenarios of ubiquitous xG networks.
+However, existing literature lacks a universal CT architecture to address the challenges of heterogeneous scenarios, data, and resources in xG networks, which hinders the widespread deployment and applications of CT.
+This article discusses a new modularized CT architecture to bridge scene recognition, cooperative sensing, and decentralized training, comprising versatile model configuration, multimodal cooperative sensing, and lightweight twin modeling modules.
+Additionally, this article presents a detailed concept, technical features, and case studies of CT, outlines mainstream trends of realization methods, followed by potential applications of CT-empowered ubiquitous connectivity, and issues requiring future investigations.
 Index Terms:
@@
 To have ubiquitous connectivity, advanced coverage extension solutions such as space-air-ground-sea-integrated network (SAGSIN), cell-free network, and reconfigurable intelligent surface (RIS) have been proposed [2].
-The provisioning of zero coverage-hole networks has to deal with the challenges related to diversified wireless channels across heterogeneous networks and multiple frequency bands, thus rendering network planning and optimization more convoluted than ever.
-This necessitates attention to one of the fundamental challenges in wireless communication systems, namely, channel acquisition, which determines the practical performance limits of wireless systems and networks.
-I-A Paradigm Shifts in Channel Acquisition
-Recently, channel twinning (CT) has emerged as a new regime of channel acquisition that reflects the complex interactions of channels and physical environments.
-CT has received increasing attention because it subverts traditional channel acquisition in the following research paradigms.
-I-A1 A shift of perspective from mathematical to environmental modeling
-Mathematical models (statistical and deterministic models) cannot balance the accuracy and complexity [3].
-Instead, CT aims to reveal the channel mapping from environmental attributes, and achieve an accurate channel acquisition with low complexity.
-I-A2 A shift from unimodal to multimodal sensing
-Traditional channel acquisition primarily relies on radio frequency (RF) sensing-based pilot training methods, accompanied by high overhead.
-In contrast, multimodal fusion is a pillar of CT, and it is promising to reduce or even eliminate RF pilots [4].
-I-A3 A shift from physical-world to digital-world measurement
-In xG networks, cross-domain and high-mobility environments can cause poor channel estimation accuracy.
-While channel prediction can apply to dynamic scenarios, it is often not environment-aware and is limited by the correlation model.
-By emulating the interactions between the channel and the physical world, CT can create real-time channel replicas in the digital world.
+The provisioning of zero coverage-hole networks has to deal with challenges related to diversified wireless channels across heterogeneous networks and multiple frequency bands, thus rendering network planning and optimization more convoluted than ever.
+This necessitates attention to one of the fundamental challenges in wireless communication systems: channel acquisition (CA), which determines the performance limits of wireless networks.
+I-A Advances in Channel Acquisition
+Recently, channel twinning (CT) has emerged as an innovative approach to CA that captures the intricate interactions between channels and physical environments.
+CT has received increasing attention for its ability to enhance traditional CA in the following ways:
+I-A1 From mathematical to environmental modeling
+It can be difficult for mathematical models (statistical and deterministic models) to strike a balance between channel prediction accuracy and computational complexity [3].
+Instead, CT aims to reveal the channel mapping from environmental attributes and achieve an accurate CA with low complexity.
+I-A2 From unimodal to multimodal sensing
+Traditional CA primarily relies on radio frequency (RF) sensing-based pilot training methods, accompanied by high overhead.
+In contrast, multimodal fusion [4, 5] is a pillar of CT, and it is promising to reduce or even eliminate RF pilots.
+I-A3 From physical-world to enhanced digital-world processing
+Traditional channel modeling or channel estimation methods often rely on channel assumptions to fit the channel model solely based on measurements.
+CT exploits measurements to extract the environment semantics or interpolate the channel measurements at any location, leading to a channel knowledge model that better reflects real-world environments.
+On the other hand, CT has the potential to significantly reduce the dependency on expensive real-world measurements once the CT model is constructed.
 Table I provides a conceptual comparison to clarify the relationship between CT and other related concepts.
 TABLE I: Comparison between different concepts about channel acquisition.
-Concept
-Role
-Execution
-phase
+{tblr}
+width = colspec = Q[120]Q[140]Q[160]Q[160]Q[120]Q[180], cell15 = c=20.3, cell21 = r=3, cell25 = c=2,r=30.3, cell43 = c=20.3, cell55 = c=20.3, cell65 = c=20.3, cell71 = r=3, cell75 = r=3, vlines, hline1-2,5-7,10 = -, hline3-4 = 2-4, hline8-9 = 2-4,6, Concept & Modeling Method Pros Cons Role and Application
+Chanel
 modeling
-Method
-Accuracy
-Generalizability
-Complexity
-Environment
-synchronization
-Research
-perspective
-Chanel modeling Derive proper models for effective channel representation Before network deployment
-Statistical
-modeling
-Low $\times$× Low $\times$× Mathematical representation
-Deterministic
-modeling
-High $\times$× High $\times$×
-Hybrid
-modeling
-Moderate $\times$× Moderate $\times$×
+Statistical Versatility Limited accuracy • Role: Derive effective model representation
+• Application: Pre-deployment
+Deterministic High accuracy Time-consuming
+Hybrid Balance accuracy and complexity
 Channel
 estimation
-Obtain current
-channel parameters
-After network
-deployment
-Pilot-based or
-training-based
-Moderate $\times$× Moderate $\times$×
-Mathematical
-representation
+Pilot-based Stable performance • High pilot load
+• Outdated CSI • Role: Obtain current CSI
+• Application: Post-deployment
 Channel
 prediction
-Infer future
-channel states
-After network
-deployment
-Time series
-analysis
-Moderate $\times$× Moderate $\times$×
-Mathematical
-representation
-Channel twinning Create the digital replica of the channel evolution Full life cycle Synthetic CT High $\times$× Low $\times$× Environment semantics
-Transfer CT Moderate Moderate
-Multiverse CT Multi-level High
+Time series analysis Proactive management Error accumulation • Role: Infer future channel states
+• Application: Post-deployment
+Channel
+twinning
+Spatial interpolation • Simple implementation
+• Low data dependency • Limited accuracy
+• Poor scalability • Role: Create the digital replica of the channel evolution
+• Application: Full life cycle Stable scenarios with smooth channel changes
+Environment semantics Enhanced modeling accuracy Expensive environment data collection Scenarios with significant environment structure effects
+Deep learning • Automatic learning
+• No explicit assumptions • Need large data
+• Hyper-parameter tuning Highly dynamic and complex scenarios
 I-B Current Status of CT Studies
-In the context of CT, some initial studies have been reported in [5, 6, 7, 8].
-The term of DT channel (DTC) [^1] To emphasize that “channel twinning” is a technology, similar to channel modeling or channel sounding, we have not used the terminology “DTC”.
+In the context of CT, some initial studies have been reported in [6, 7, 8].
+The term of digital twin channel (DTC) [^1] To emphasize that “channel twinning” is a technology, similar to channel modeling or channel sounding, we have not used the terminology “DTC”.
 appeared earlier in [6], but the concept has not been specified.
-Later, the potential applications of predicting channels via digital twin (DT) were highlighted in [6], although the DTC concept was not proposed.
-The authors of [7] mentioned that DTC should enable channel prediction but did not provide clear definitions.
+Later, the potential of channel recreation on digital twin (DT) was revealed in [7], although the DTC concept was yet to be proposed.
 More recently, DTC has been defined as the mapping between channels and associated communication operations [8].
-Therein, the feasibility of real-time DTC has been demonstrated experimentally.
-As an initial instantiation of the CT concept, Boston Twin [9] was developed to introduce a city-scale twinning system for wireless networks.
-The above literature has focused primarily on beam prediction applications or only discusses the knowledge graph construction from a local and centralized perspective.
-However, deploying CT in xG networks faces the challenges of cross-domain modeling, distributed sensing, and lightweight training.
-A centralized strategy makes it hard to cater to heterogeneous resources in xG networks.
-Therefore, in this article, a systematic discussion on the modularized architecture of ubiquitous CT is conducted to facilitate the organic integration of CT and ubiquitous sensing and intelligence.
+Therein, the radio environment knowledge pool (REKP) was proposed to support DTC.
+The above literature has focused primarily on channel knowledge map (CKM) applications [2] or only discusses the REKP construction from a local and centralized perspective.
+While providing a framework for reflecting propagation environment via channel fingerprints, CKM emphasizes static mappings.
+REKP [8] is a knowledge semantic database that stores and manages the feature relationships of radio environments and channels.
+However, it is not tailored to specific scenarios or environments.
+In xG networks, deploying CT faces the challenges of cross-domain modeling, distributed sensing, and lightweight training.
+A centralized strategy makes it hard to cater to heterogeneous resources.
+In this article, a systematic discussion on the modularized architecture of ubiquitous CT is conducted to facilitate the organic integration of CT and ubiquitous sensing and intelligence.
 For a better understanding of CT, a concise definition and several clarifications regarding the general CT misconceptions are refined.
 This article also highlights the technical evolutions, the potential use cases for CT-empowered ubiquitous connectivity technologies, and several promising directions to spur future research.
-Fig.
-1: Conceptual framework of CT in xG networks with ubiquitous connectivity.
-Advanced ubiquitous connectivity scenarios give birth to diversified wireless channels across heterogeneous networks.
-All of these channels can be acquired in the CT system without the need for high-overhead channel estimation.
 II Basic Concepts and Features of CT
 II-A Concepts and Realization Principle of CT
-In this article, CT is defined as a digital mirror construction of physical channels for network design, optimization, and evaluation without real deployment.
-Essentially, such a digital mirror is a mapping model that indicates the correspondence from the environment to the associated channel state information (CSI).
-Specifically, a site-specific CT attempts to find a mapper to approximate the RF propagation law.
+CT is defined as a digital mirror construction of physical channels for network design, optimization, and evaluation without real deployment.
+In essence, such a digital mirror is a mapping model that indicates the correspondence from the environment to various aspects of associated channel behaviors, such as channel state information (CSI), channel path angle, or channel gain, depending on the specific task.
+Specifically, a site-specific CT aims to simulate the RF propagation law.
 As shown in Fig.
-1, given the environmental conditions, CT can acquire the CSI at any position.
-Regarding the realization of CT, current mainstream methods can be primarily divided into three categories: Spatial interpolation modeling, environment semantic modeling, and deep learning (DL) modeling.
-II-A1 Spatial Interpolation Modeling
-Spatial interpolation [2] only utilizes channel measurements within the region of interest (RoI) to predict the channel of any location.
-II-A2 Environment Semantic Modeling
-Environment semantic-based methods, such as spatial loss field and virtual obstacle model [10], estimate the fading distribution of the entire RoI by observations.
-II-A3 Deep Learning Modeling
+1, CT can acquire CSI at any position based on the environmental conditions.
+Fig.
+1: Conceptual framework of CT in xG networks and its modularized CT architecture.
+The versatile model configuration module provides information for measurements and model design to the other two modules.
+It generates synthetic data to pretrain the CT model in the lightweight twin module.
+Then, the CT model can be fine-tuned by feeding data from the multimodal cooperative sensing module.
+Current methods for realizing CT fall into three main categories: Spatial interpolation (SI) modeling, environment semantics (ES) modeling, and deep learning (DL) modeling.
+II-A1 SI
+SI leverages the spatial correlations of sampling channel measurements within the region of interest (RoI) to predict the channel of unobserved locations.
+Classic SI methods include inverse distance weighting, Kriging, etc., commonly used for constructing CKM [2].
+II-A2 ES
+Utilizing ES features (e.g., building layout, terrain type, and obstacles), ES modeling can help accurately build channel models, such as the virtual obstacle model [9] and REKP [8].
+II-A3 DL
+DL methods apply advanced deep networks to learn the complex mapping between various environmental factors affecting the channels.
 In DL methods, differentiable ray tracing (DRT) is gaining popularity as a promising DT enabler [10].
@@
 In practice, suitable methods should be selected according to applications and dataset sizes.
-From the perspective of implementation complexity, CT deployment should undergo three evolution stages: i) Synthetic CT [7], which is deployed in a quasi-static environment to generate synthetic data that approximates the measurements in the RoI; ii) Transfer CT [12], which constructs self-adaptive CT based on knowledge transfer to cope with unforeseen scenarios; and iii) Multiverse CT [13], which customizes different twins for applications with varying computing power and delay constraints.
-Each twin model offers multiple fidelity levels with different processing delays.
-II-B Common Misconceptions and Three Clarifications
-To elaborate on the concept of CT comprehensively, we discuss the CT model illustrated in Fig.
+From the perspective of implementation complexity, CT deployment should undergo three evolution stages: i) Synthetic CT [7], which is deployed in a quasi-static environment to generate synthetic data that approximates the measurements in the RoI; ii) Transfer CT [11], which constructs self-adaptive CT based on knowledge transfer to cope with new scenarios; and iii) Multiverse CT [12], which customizes different twins to offer multiple fidelity levels for applications with diverse computing power and delay constraints.
+II-B Common Misconceptions and Four Clarifications
+To elaborate on the concept of CT, we discuss the CT model illustrated in Fig.
 1 from inputs to outputs.
-Several basic features are highlighted as follows: a) context-awareness: Instead of merely inputting received pilots or feedback by RF sensing methods, CT integrates contextual information from multimodal sensors (e.g., camera, GPS and Radar) to promptly adapt to environmental dynamics; b) high-fidelity: CT allows for high-fidelity emulation of complex interactions between the environment and the network.
-Thanks to three-dimensional (3D) map-aided methods, e.g., ray-tracing, spatial consistency [1] and multi-hop propagation can be simulated; c) self-synchronization: CT can model the whole life cycle of channel behaviors in the wireless environment for real-time emulation; d) expertise-driven: EM and channel modeling expertise are needed to design CT models.
-This helps produce compact DL structures and improve interpretability when specific channel characteristics, e.g., signal-to-noise ratio (SNR), are desired to be output; and e) region-specific: Since the CT model is trained on measurements from a specific site’s coverage region, the DTC must be region-dependent.
-Since CT is an emerging technology, there are several common misconceptions that need to be clarified.
-Clarification 1: Do widely-used statistical channel modeling techniques belong to CT technology?
-Following the CT concept, a DTC refers to a digital mirror of the channel response in a given environment, whereas standard statistical models like Rayleigh and Rician, cannot match specific environments.
-Moreover, 3GPP urban microcell (UMi) and urban macrocell (UMa) models only consider macroscopic environmental attributes (e.g., urban or rural) but ignore the specific environmental factors.
-In this sense, statistical models may not create the real DTCs.
+Several basic features are highlighted as follows: a) context-awareness: Instead of merely using received pilots or feedback, CT integrates contextual information from multimodal sensors (e.g., camera, GPS, and Radar) to learn environmental dynamics; b) high-fidelity: Thanks to three-dimensional (3D) map-aided methods [9, 10], e.g., ray-tracing, CT allows for high-fidelity emulation of complex interactions between the environment and the network; c) self-synchronization: CT can model the whole life cycle of channel behaviors in the environment for real-time emulation; d) expertise-driven: EM and channel modeling expertise help design compact DL structures and improve interpretability when specific channel properties, e.g., signal-to-noise ratio (SNR), are desired to be output; and e) region-specific: Since the CT model is trained on measurements from a specific site’s coverage region, the DTC must be region-dependent.
+Since CT is an emerging technology, several common misconceptions need to be clarified.
+Clarification 1: Do statistical and deterministic modeling belong to CT technology?
+CT is designed to create a digital mirror of the channel response in a given environment, whereas standard statistical models like Rayleigh and Rician cannot match specific environments.
+Besides, 3GPP urban microcell (UMi) and urban macrocell (UMa) models only consider macroscopic environmental attributes (e.g., urban or rural) but ignore the specific environmental factors [3].
+In this sense, statistical models may not reflect the CT features.
+Deterministic models, such as ray tracing, is a physics-based model that predicts channel by simulating ray propagation in specific environments and scene geometries, as well as complex EM interactions with surrounding objects.
+On the other hand, CT is a mapping model-based CA method that seeks to efficiently mimic channel behaviors with no need for exhaustive recursive tracing.
+As a preliminary attempt at CT, Sionna is ten times faster than traditional ray-tracing like Wireless InSite (WI) [10].
+Conventional channel modeling methods lay a solid foundation for understanding the physical behaviors of waves, while CT can adaptively respond to environmental changes through its self-synchronization capability, as discussed in Section II-B.
+Combining ray tracing with CT can further contribute to advanced channel modeling, and potentially bring real-time ray tracing to reality.
 Clarification 2: Does CT necessarily mean the use of DL?
-According to specific application requirements, one can flexibly choose different CT training methods, e.g., spatial interpolation, environment semantic, and DL methods.
+According to specific application requirements, one can flexibly choose different CT training methods, e.g., SI, ES, and DL modeling methods.
 Clarification 3: Does using only position information to predict network operations belong to CT technology?
 CT is either explicitly or implicitly context-aware.
-Only position-based methods construct the channels under the simple line-of-sight (LoS) assumption that ignores the environmental factors.
+Only position-based methods construct the channels under the simple line-of-sight (LoS) assumption.
 However, blockages and scatters may cause non-LoS (NLoS) cases.
 This shows the significance of context awareness, and that position-based channel recovery cannot compete with CT.
+Clarification 4: What are the conceptual differences between DT, CT, and DTC?
+A DT is a virtual replica of a physical entity (e.g., a network or device).
+DTC develops an REKP to help understand the interaction between channels and environmental characteristics across extensive scenarios [8].
+The REKP framework updates the path loss feature relationship graphs for different receiver locations through real-time acquisition of environmental features (e.g., position, blockage, and distance).
+However, these updates rely on static environmental factors within localized time intervals, rather than comprehensive dynamic modeling of channel evolution over time.
+In contrast, CT focuses on learning both static attributes and the temporal evolution of channels, emphasizing the twinning process in a specific region.
 III A Modularized Design of CT Architecture
@@
   • Q1.
-    Since xG networks are anticipated to involve massive cross-domain links across space/frequency/time-domains, how can we make the channel modeling of CT applicable to any scenario?
+    Since xG networks are anticipated to involve massive cross-domain links across space/frequency/time domains, how can we streamline the CT model deployment and configuration process for diverse scenarios and channel types?
   • Q2.
@@
 In light of these questions, we present a modularized architecture to discuss the design considerations for CT deployment in xG networks.
-The overall CT architecture comprises three modules of universal channel modeling, multimodal cooperative sensing, and lightweight decentralized CT training, as shown in Fig.
-2.
-Fig.
-2: The overall modularized CT architecture.
-The universal channel modeling module provides information about the sensing targets and prediction tasks of interest for the other two modules.
-The channel modeling module can generate synthetic data, which can be leveraged to pretrain the CT model in the lightweight twin module.
-Then, the CT model can be fine-tuned by feeding sensing measurements from the multimodal cooperative sensing module.
-III-A Universal Channel Modeling Module
+The overall CT architecture comprises three modules of versatile model configuration, multimodal cooperative sensing, and lightweight decentralized CT training, as shown in Fig.
+1.
+III-A Versatile Model Configuration Module
 Various scenarios in xG networks give rise to various types of channels.
 The characteristics of these channels differ significantly from each other.
-Consequently, a major challenge that arises for CT deployment is how to integrate the diverse channel characteristics across scenarios into a unified model.
-To this end, a universal channel modeling module tailored for ubiquitous CT is needed.
+Consequently, a major challenge that arises for CT deployment is how to integrate the diverse channel characteristics across scenarios to facilitate the automated construction of suitable CT models.
+To this end, a versatile model configuration module tailored for ubiquitous CT is needed.
 Generally, a channel consists of large-scale fading (LSF) and small-scale fading (SSF) components, each with different characteristics in various scenarios.
-Therefore, the channel components associated with the scenario and working band should be identified at first.
-For instance, high-mobility scenarios, such as Internet of Vehicles (IoVs) and unmanned aerial vehicle (UAV) networks, should consider non-stationary channels.
+Therefore, the channel components associated with the scenario and working band should be identified first.
+For instance, high-mobility scenarios, such as Internet of Vehicles (IoV) and unmanned aerial vehicle (UAV) networks, should consider non-stationary channels.
 Gas absorption and blocking effects should be considered in millimeter-wave (mmWave) and Terahertz (THz) networks, while visible light communication (VLC) channels only involve LSF.
@@
 To effectively deploy CT, we can quickly determine the channel components by following two steps: 1) LSF channel construction by a static 3D map; and 2) SSF channel prediction by capturing dynamic activities.
+The versatile model configuration module allows developers to easily configure the CT model by selecting from various model settings and adjusting parameters to suit specific task needs.
+Additionally, this module provides supplementary synthetic data to assist with CT model training, rather than engaging in traditional channel modeling.
 III-B Multimodal Cooperative Sensing Module
@@
 In the multimodal cooperative sensing module shown in Fig.
-2, there are two main categories of multimodal sensors: RF and non-RF.
-The traditional channel measurement sources come primarily from RF reception and sensing.
+1, there are two main categories of multimodal sensors: RF and non-RF.
+The traditional channel measurements come primarily from RF reception and sensing.
 Such measurement is limited by the sensing range and accuracy of the single-site ISAC scheme.
-For RF sensing, a multi-site cooperative sensing scheme can be preferable, which can enhance signal acquisition by virtue of ubiquitous connectivity.
-In this scheme, non-terrestrial and terrestrial base stations (BSs) can sense any position through ISAC technology.
+For RF sensing, a multi-site cooperative sensing scheme is preferable, which can enhance signal acquisition by ubiquitous connectivity.
+In this scheme, non-terrestrial and terrestrial base stations (BSs) can sense any position through ISAC.
 In addition, users can also track targets by the dedicated positioning reference signals defined in 5G-NR R17 [13].
@@
 In this sense, the challenge of M2SDF lies in how to deal with the data heterogeneity at the CT server.
-To extract propagation environment semantics, semantic alignment and completion [8] techniques may be invoked.
+ES plays a crucial role in enabling CT by providing detailed information about the physical environment.
+The ES implementation in the CT framework consists of the following steps: 1) Extracting physical structure features of the environment that impact signal paths, such as walls, trees, and other obstructions [9]; 2) classifying and encoding these environment features to construct channel semantic information.
+To extract propagation ESs, semantic alignment, and completion [8] techniques may be invoked; and 3) utilizing semantic features to model the interactions between network actions and specific tasks.
+Thus, optimization algorithms or deep networks can be designed to improve the accuracy of CA, adapt to dynamic scenarios, or enhance communication performance.
 Fig.
-3: Training architecture of CT: (a) Centralized learning-based CT model construction, where dedicated measurement anchors collect data and upload data to the CT server for training; and (b) FL-based CT model construction, where registered device clients use their local data to train the model and only upload trained parameters for model aggregation at the CT server.
+2: Training architecture of CT: (a) Centralized learning-based CT, where dedicated measurement anchors collect data and upload it to the CT server for training; and (b) FL-based CT, where registered device clients use their local data to train the model and only upload trained parameters for model aggregation at the CT server.
 III-C Lightweight Twin Modeling Module
 Existing CT models typically train in a centralized manner.
-However, scalability and privacy challenges limits the CT deployment in ubiquitous connectivity.
+However, scalability and privacy challenges limit the CT deployment in ubiquitous connectivity.
 In this case, federated learning (FL) is viewed as an adequate and promising choice for lightweight CT because FL training is decentralized and does not require sharing data.
@@
 2 shows a lightweight CT training module based on FL.
-In the FL-enabled CT scheme, FL clients only need to upload local model parameters to collaboratively train a global CT model, thus reducing communication costs and privacy leakage risks.
+In the FL-enabled CT scheme, FL clients only need to upload local model parameters to train a global CT model collaboratively, thus reducing communication costs and privacy leakage risks.
 The integration of FL and DT is not straightforward, as two practical issues need to be addressed as prerequisites.
-One is that participating in FL tasks can increase the energy costs of client devices.
-For this reason, self-interested sensor devices may be reluctant to participate in FL, which needs proper FL incentive mechanisms [15].
-Clients participating in FL should meet their own energy constraints.
-For example, UAVs are energy-limited.
-In this sense, an incentive mechanism for FL-based CT should be designed to improve CT training efficiency while meeting the power constraints.
-To achieve such energy-efficient FL incentives, we need to quantify the contribution of clients to the global CT model convergence and design appropriate payment mechanisms.
-The other practical issue that needs to be addressed for the integration of FL and DT is the heterogeneity of local data on each client.
-This would require handling data distribution imbalances across clients.
-The reason is that personalized FL may be suitable for lightweight CT in the future.
+First, participating in FL tasks can increase the energy costs of client devices.
+Self-interested devices may be reluctant to participate in FL.
+Thus, proper FL incentive mechanisms [14] are needed to encourage participation while considering energy constraints, especially for energy-limited devices like UAVs.
+To achieve such energy-efficient FL incentives, we need to quantify the contribution of clients to the global CT model convergence and design suitable payment mechanisms.
+Second, the heterogeneity of local data across clients requires handling data distribution imbalances, as personalized FL could become promising for lightweight CT in the future.
 IV Case Studies: Intelligent Coverage and Lightweight Twin
 This section provides two case studies to demonstrate the significance of CT for ubiquitous connectivity and discuss the lightweight FL-based CT performance.
+Fig. 3: User association results produced based on the ground-truth channel, the twin channel, and the fitted PL model.
+IV-A CT-Based Intelligent Coverage
+In ubiquitous connectivity networks, user association can dynamically adjust the connection of user equipment (UE) according to channel conditions to ensure continuous and efficient coverage.
+By selecting optimal access points (APs) for UE access, user association can improve resource utilization, reduce packet loss, and prevent interruptions.
+However, accurate full CSI is essential for effective user association.
+Cell-free communication is taken as an example to illustrate the benefits of CT for ubiquitous connectivity.
+In cell-free networks, obtaining the full CSI of involved nodes is challenged by limited pilots and high CSI exchange overhead.
+To address this, we propose a CT-based intelligent user association scheme.
+Specifically, the scheme is implemented via offline CT model training and online CSI query, as illustrated in Fig.
+2(a).
+During the offline phase, dedicated measurement anchors collect channel measurements in the RoI and upload measurements to the CT server for training.
+Once trained, the CT model is stored in the twin repository.
+During the online phase, UEs request the CT server to query the CSI of the node they attempt to communicate with.
+We leverage the synthetic data generated by the WI software [8] to train a CT model.
+As shown in Fig.
+3, the transceiver locations are selected from grid points with a uniform spacing of 8 meters.
+We then use a deep neural network with seven hidden layers to capture how the environment affects the channel gain distribution.
+The network takes four-dimensional transceiver coordinates as input and outputs the corresponding channel gains.
+For any given transmitter location, we perform SI [2] using the predicted values from 30 randomly selected receiver positions, enabling the generation of the channel gain maps.
 Fig.
-4: User association results produced based on the true channel, the twin channel, and the fitted PL model: (a) True channel gain; (b) twin channel gain; (c) channel gain predicted by fitted PL model; and (d) the achievable sum rate with different schemes.
-Each UE selects five APs for communications, and the associated APs are indicated by red dots.
-Fig.
-5: Federated twin of the channel gains: (a) The validation loss versus communication rounds, where the total number of training samples is 10,000, and these samples are randomly divided into three or five UEs for FL; and (b) MAE comparison of channel gain prediction across different models.
-IV-A CT-Based Intelligent Coverage
-In ubiquitous connectivity networks, user association can dynamically adjust the connection of user equipment (UE) according to the channel status, in order to achieve continuous and efficient coverage.
-By user association, the optimal access points (APs) are selected for UE access, affecting network resource utilization.
-Dynamic adaptive user association can reduce latency or packet loss, avoid interruption, and save the transmit powers.
-However, the premise of dynamic adaptive user association is an accurate perception of full CSI.
-Cell-free communication, as a key ubiquitous connectivity technology, is taken as an example to illustrate the advantage of CT.
-In cell-free networks, obtaining the full CSI of involved nodes is challenged by limited pilots and high CSI exchange overhead.
-To address this, we propose a CT-based intelligent user association scheme in cell-free networks.
-Specifically, the proposed scheme is implemented via offline CT model training and online CSI query, as illustrated in Fig.
-3(a).
-In the offline phase, dedicated measurement anchors are responsible for collecting channel measurements at various locations in the RoI and uploading measurements to the CT server for training.
-Once the CT model is obtained through offline training, it is stored in the twin repository.
-In the online phase, APs or UEs request the CT server to query the channel information of the node they attempt to communicate with.
-To demonstrate the advantage of CT, we leverage the synthetic data generated by Wireless InSite [^2] https://www.remcom.com/wireless-insite-propagation-software software for model training [2].
-Channel gain measurements, sampled at a uniform grid in the region shown in Fig.
-4(a), serve as the ground truth.
-Based on the spatial interpolation concept, the trained CT model can capture the impact of environmental factors on channel gains.
-Fig.
-4(b) shows a channel gain map given by a UE position, where the UE is indicated by a white square.
-Conventional user association is based only on the distance or path loss (PL) in practical cell-free systems.
-We derive the PL model with the same synthetic data by least squares fitting.
-As a benchmark, PL-based results are shown in Fig.
-4(c).
-By comparing Figs.
-4(a), 4(b), and 4(c), it can be observed that CT can efficiently perceive the environment, while the fitted PL model cannot reproduce the environment profile.
-With the predicted channel gain, we select five APs from 20 distributed APs within the RoI for user association.
-Since the PL model does not grasp the environment semantics, it can only select APs according to the distance criterion.
-The channel gain comprises PL, shadowing, and multi-path effects.
-PL is determined by distance while shadowing and multi-path effect are affected by surroundings.
-By contrast, CT can guide the UE in avoiding buildings, instead of directly selecting the nearest APs.
-Fig.
-4(d) shows the sum rates achieved by different user association schemes, which confirms the benefits of CT for ubiquitous coverage.
-In addition, since open street maps [10] and user locations are reasonably easy to obtain, the CSI of any AP-UE link can be retrieved by CT.
-Thus, the dreadful CSI-sharing overhead can be avoided in ubiquitous connectivity.
+3 shows channel gain maps with the UE marked by a white square.
+Conventional user association relies on the distance or path loss (PL) only.
+We fit a PL model with the same synthetic data by least squares fitting as a benchmark.
+By comparing the maps in Fig.
+3, CT can efficiently perceive the environment, while the fitted PL model fails to reproduce the environment profile.
+With the predicted channel gain, we select five out of 20 distributed APs within the RoI for user association.
+The PL model can only select APs according to the distance criterion, as it does not grasp the ESs.
+By contrast, CT helps the UE avoid obstacles rather than selecting the nearest APs.
+In addition, OpenStreetMap (OSM) data [9] can assist the CT model by generating synthetic data, particularly in scenarios where RF measurements might be sparse or unavailable, the CSI of any AP-UE link can be retrieved by CT.
+Fig. 4: Validation loss of MSE versus communication rounds, where the total number of training samples is 10,000, and these samples are randomly divided between three or five UEs for FL.
+Fig. 5: MAE comparison of channel gain prediction across different models.
 IV-B Federated Learning-Based Lightweight CT
@@
 In light of this, we investigate an FL-based lightweight channel gain twinning model; see Fig.
-3(b).
-In the FL-based CT training architecture, no dedicated measurement anchors are needed for data collection.
-Each mobile UE may be a potential measurement anchor.
-These incentive UEs can record their locations and associated channel measurements as they move into the RoI.
-Unlike centralized training, the UEs do not upload the data; rather, they first register their identities with the CT server’s client manager for subsequent FL training.
+2(b).
+In the FL-based CT training architecture, no dedicated measurement anchors are needed.
+Mobile UEs act as potential measurement anchors.
+Incentive UEs can record their locations and associated channel measurements as they move into the RoI.
+Unlike centralized training, the UEs first register their identities with the CT server’s client manager for subsequent FL training.
 When the CT server prepares to train a CT model, a group of registered UEs train their own CT models locally by using their recorded measurements.
-Then, they upload only the trained parameters to the server for model aggregation, thus obtaining the global CT model.
-The CT server sends the global model parameters to the registered clients to update the local models.
+Then, they upload only the trained parameters to the server for model aggregation.
+The CT server updates the global model and sends model parameters back to the registered clients for local model updates.
 After several rounds of interactions, the global model converges, and hence, the FL-based CT training is completed.
-We select a portion of the channel gain dataset used in the previous subsection, randomly splitting and distributing it to FL clients.
+A portion of the channel gain dataset was randomly split among FL clients.
 Fig.
-5(a) shows the trained results.
-It is seen that more clients contribute to better validation loss performance, even if the total data samples remain the same.
-With more clients, the gradient noise of individual clients can be smoothed out.
-This indicates that increasing the number of clients can improve the federated twin performance, but raises communication overhead concerns.
+4 shows the training loss of mean squared error (MSE).
+It is seen that more clients contribute to lower MSE loss, even with the same total data samples.
+This is because more clients help smooth out the gradient noise of individual clients and improve the federated twin performance, though communication overhead rises.
 Fig.
-5(b) compares prediction performance among the centralized CT model, fitted PL model, and federated CT model in terms of mean absolute error (MAE).
-From the boxplots of MAE, the red solid line denotes the median MAE, the box represents the distribution range of the majority of MAE, and the black dots outside the box denote outliers.
-It is observed that the PL model exhibits high prediction errors overall, indicating poor prediction accuracy.
-Meanwhile, few outliers indicate prediction stability, as the channel gain produced by the PL model lacks shadowing and SSF compared to the true channel.
+5 compares the channel gain (dB) prediction capabilities in terms of mean absolute error (MAE).
+The yellow solid line denotes the median MAE, the box represents the distribution range of the majority of MAE, and the black dots outside the box denote outliers.
+The PL model exhibits high prediction errors overall, indicating poor prediction accuracy.
+Meanwhile, few outliers indicate prediction stability, as the channel gain produced by the PL model lacks shadowing compared to the real channel.
 Both CT models show better prediction accuracy.
-Even decentralized training-based CT model has superior accuracy to the PL model.
 Despite the fluctuations in prediction error of CT models, occasional outliers do not affect the overall performance of ubiquitous connectivity.
-Notably, the goal of CT is not to copy physical channel entities, but rather to satisfy network requirements by emulation.
+In FL, the primary feedback overhead arises from the transmission of model updates from the UE to the BS.
+The feedback overhead of the 3GPP Type II codebook scales with the rank, and requires hundreds of bits (more than 500 bits for some system setups) [15].
+Since the overhead of different setups varies greatly, it is difficult to pre-determine the required overhead.
+Thus, the BS allocates physical uplink shared channel (PUSCH) resources for CSI reporting according to the maximum overhead, which causes resource waste and reduced efficiency.
+In this case, the UE can utilize leftover PUSCH resources for FL model parameter uploads after completing CSI reporting, thus improving resource utilization.
+Additionally, once the FL model is trained for large time-scale invariant environments, it will significantly reduce the channel feedback overhead.
 V Convergence of CT and Ubiquitous Communications: Opportunities and Challenges
+In addition to reducing pilot overhead and enhancing prediction accuracy, CT can benefit ubiquitous connectivity by offering enhanced dynamic system robustness [12], accelerated network development and testing [3], and improved resource management [4].
+To unfold the potential of CT, this section presents some open issues concerning the convergence of CT and ubiquitous communications.
 V-A Opportunities of CT for Ubiquitous Connectivity
@@
 Due to the down-tilted antenna pattern, UAVs may easily fall into the antenna nulls of nearby BSs.
-Therefore, the UAVs must search for the best antenna sidelobe associated with the remote cell when switching cells instead of using the aforementioned distance criterion.
-In this case, CT can emulate the antenna gain and airframe shadowing [1] to achieve high-efficiency cell association in UAV networks.
+Therefore, the UAVs cannot use the distance criterion to choose the cell.
+In this case, CT can emulate the antenna gain and airframe shadowing [1, 4] to achieve high-efficiency cell association in UAV networks.
 V-A2 CT-Empowered IoV Communications
-One of the most immediate applications of CT is predictive beamforming in IoVs.
-CT-enabled beamforming must only send necessary control signaling without extra pilots [4].
-On the other hand, RIS is regarded as an attractive low-power technology that can overcome dynamic blockage in IoVs.
+One of the most immediate applications of CT is predictive beamforming in IoV.
+CT-enabled beamforming requires only necessary control signaling without extra pilots [5].
+On the other hand, RIS is regarded as a low-power solution to dynamic blockage issues in IoV.
 However, CSI acquisition and frequent beam alignment in high-mobility scenarios hinder the widespread application of RIS.
@@
 V-A3 CT-Empowered Maritime Communications
-CT is promising in achieving ubiquitous maritime connectivity.
-Note that the maritime communications environment has distinct features of sparse scatters compared to terrestrial networks.
-This motivates us to employ the environmental information to deploy CT models for network planning [3].
+Since the maritime communications environment has distinct features of sparse scatters compared to terrestrial networks, CT is expected to enhance maritime connectivity by employing environmental information for network planning [3].
 Channel interference can be acquired in advance to avoid the acquisition burden of full CSI.
 In the harsh underwater environment, underwater acoustic (UWA) channels suffer from spatio-temporal-frequency variability, limited bandwidth and severe noise and interference [1].
-Channel estimation and data transmission becomes even more challenging in multi-node UWA communications.
+Channel estimation and data transmission become even more challenging in multi-node UWA communications.
 In this case, CT also brings unprecedented opportunities for UWA communications.
@@
 V-B1 Data Deficiency
-CT demands sufficient high-quality measurement data.
-However, channel measurement campaigns are time-consuming and expensive.
-In practice, only sparse measurements from limited points are available for CT training.
-For example, DL-based indoor propagation prediction [12] only utilizes channel measurements from 36 locations.
+CT demands sufficient high-quality measurement data, but channel measurement campaigns are costly and time-consuming.
+In practice, only sparse measurements from limited points are available.
+For example, DL-based indoor propagation prediction [11] only uses measurements from 36 locations.
 Furthermore, some measurements may contain errors due to hardware defects.
-Therefore, for high-precision CT training, it is necessary to study the efficient utilization of synthetic and measurement data.
+To achieve high-precision CT, it is necessary to study the efficient utilization of synthetic and measurement data.
 V-B2 Personalized FL Mechanism of CT
 Since FL clients only collect channel measurements from the local irregular areas, a simple federated averaging-based aggregation strategy can hardly be optimal.
-In particular, personalized FL with incentive mechanisms needs to consider the CT model contribution of each client to the entire environment map.
+In particular, personalized FL with incentive mechanisms [14] needs to consider the CT model contribution of each client to the entire environment map.
 V-B3 Malicious CT Deployment
 Malicious nodes with CT ability may exist in networks.
-They can train the CT models based on local radio measurements or query the CT server for CSI at locations of interest.
-Therefore, once a malicious CT is able to recover CSI, wireless networks will face unparalleled security threats.
+They may access CSI at any location of interest.
+Therefore, once a malicious CT can recover CSI, wireless networks will face unparalleled security threats.
 Existing physical layer security and secret key schemes will be rendered ineffective.
@@
 While DTC is effective for wireless network optimization, high-dimensional channel matrices caused by large-scale MIMO or RIS raise computational and caching challenges for CT deployment.
-In this case, an equivalent channel compression representation scheme in CT is attractive for network optimization.
+In this case, an equivalent channel compression representation [7] scheme in CT is attractive for network optimization.
 VI Conclusion
-From the perspective of the physical environment, a new channel acquisition paradigm, namely, CT technology, has been studied to support xG ubiquitous wireless communication networks.
+From the perspective of the physical environment, a new CA paradigm, CT technology, has been studied to support xG ubiquitous wireless communication networks.
 With multimodal sensing, CT can evolve to predict CSI that closely resembles actual channels.
@@
 To promote CT deployment in wireless networks, a modularized design architecture of CT has been presented.
-The architecture includes universal channel modeling of full scenarios, multimodal cooperative sensing, and FL-based lightweight CT training.
+The architecture includes versatile model configuration of full scenarios, multimodal cooperative sensing, and FL-based lightweight CT training.
 New opportunities for CT in enabling technologies for ubiquitous connectivity and future research challenges have also been discussed.